A public university hires students to solicit donations by telephone. After a brief training period, the students call alumni and are paid on a commission basis. Experience indicates that early on, these students tend to have only modest success and that 65% of them quit their jobs in their first two weeks of employment. The group hires six students which can be viewed as a random sample. What is the probability that less than half of the new hires quit within the first two weeks?
Answer #1

This will be the binomial distribution with parameters:

n = 6, p = 0.65

Hence,

P(Less than half)

= P(X < 3)

= P(X = 0) + P(X = 1) + P(X = 2)

= 0.1174